OpenStudy (anonymous):

Solve for x (3x)^6=3^-6

OpenStudy (anonymous):

We can take the cube root of both sides to get (3x)^2 = 3^(-2) Now we can "take the square root", but must keep in mind that there is a positive and a negative answer, so that \[3x = \pm(3^{-1}) = \pm{1 \over 3}\]and hence \[x = \pm{1 \over 9}\]I'm assuming x is only allowed to be real. There are four additional complex solutions to this (which would come out at the point that I took the cube root), but I won't bother typing it out unless you want/need to know what they are.

OpenStudy (anonymous):

the answer is 1/9 ?

OpenStudy (mertsj):

Write both sides with the exponent of 6

OpenStudy (anonymous):

positive or negative 1/9 Because a negative to an even power is positive.

OpenStudy (mertsj):

\[(3x)^6=(\frac{1}{3})^6\]

OpenStudy (anonymous):

no it (3^x)^6=3^-6

OpenStudy (mertsj):

Since the expressions are equal and the exponents are the same, the bases must also be the same so write: \[3x=\frac{1}{3}\]

OpenStudy (mertsj):

Solve that.

OpenStudy (mertsj):

\[(3^x)^6=3^{6x}\]

OpenStudy (mertsj):

So: \[3^{6x}=3^{-6}\]

OpenStudy (mertsj):

So 6x=-6

OpenStudy (mertsj):

x=-1

OpenStudy (anonymous):

thank you !!!!!!!!!!!!!!!!!!!!

OpenStudy (mertsj):

yw